# AeroSketch: Near-Optimal Time Matrix Sketch Framework for Persistent, Sliding Window, and Distributed Streams

### v1

@@ 1. Introduction
Many types of real-world data—such as computer network traffic (Xia et al., 2014; Kim et al., 2020), online machine learning (Kuzborskij et al., 2019; Chen et al., 2020; Wen et al., 2025; Chen et al., 2022), and sensor data (Xie et al., 2023; Aquino et al., 2007)—are continuously generated as vast, high-velocity streams of vectors (Zhou, 2024; Wang et al., 2024; Muthukrishnan, 2005). Storing even a reasonable portion of these streams in memory is impractical. Therefore, it is necessary to design online algorithms that process one vector at a time and immediately discard it, enabling real-time processing with reduced memory usage at the cost of some precision. Such compact representations are known as matrix sketching (Cormode and Yi, 2020; Liberty, 2013; Woodruff, 2014; Desai et al., 2016; Kacham, 2024).
@@ 2.1.2. Matrix Sketching over Sliding Window
Besides the space-optimal Fast-DS-FD algorithm proposed by Yin et al. in (Yin et al., 2024) as we listed in Table 1, they also mentioned a time-optimized version called probabilistic Fast-DS-FD in Theorem 3.1, claiming that the per-update time complexity is $O(d\ell)$, by setting $\ell\leftarrow\min\left(\lceil\frac{1}{\varepsilon}\rceil,d\right)$, so that the time complexity becomes $O\left(\frac{d}{\varepsilon}\right)$. Unfortunately, probabilistic Fast-DS-FD lacks a concrete implementation and therefore remains a potential direction for future work. It also lacks strict proofs of error bounds and complexities. Important parameters are not defined explicitly, such as how to set the number of iterations for the RandNLA algorithm to meet the error bound. Detailed experimental results on runtime overhead are also not provided. Therefore, we did not list it in Table 1 as a competitor to our AeroSketch framework. Additionally, we found that Lemma 1 in (Yin et al., 2024), which involves the key operation of dumping and restoring snapshots and the associated correctness proof, holds only under the premise that $\bm{v}_{j}$ is an exact singular vector. However, the probabilistic randomized algorithm used by probabilistic Fast-DS-FD to optimize time can obtain only an approximate estimate of $\bm{v}_{j}$. Therefore, directly replacing the SVD in Fast-DS-FD with RandNLA will introduce cumulative restoring norm errors, as shown in Figure 1. Designing the dumping and restoring snapshots operation and correctness proof under this premise is far from straightforward. In this paper, our AeroSketch framework will overcome these limitations.
@@ 2.1.4. Approximate Matrix Multiplication over Sliding Window
Remark 0.
Regarding Problem 4, we list two space-optimal baseline algorithms in Table 1: ML-SO-COD proposed by Xian et al. (Xian et al., 2025) and hDS-COD proposed by Yao et al. (Yao et al., 2025). Among them, (Xian et al., 2025) explicitly states in Theorem 3 that the per-update time complexity of ML-SO-COD is $O\left(\left(\left(d_{x}+d_{y}\right)\ell+\ell^{3}\right)\log R\right)$. Meanwhile, (Yao et al., 2025) claims in Theorem 1 that the per-update time complexity of hDS-COD is $O\left(\frac{d_{x}+d_{y}}{\varepsilon}\log R\right)$. However, we note that the proof of Theorem 1 writes: “the amortized time complexity for a single update step in Algorithm 4 (hDS-COD) is $O((d_{x}+d_{y})\ell+\ell^{3})$. Assuming $d_{x}+d_{y}=\Omega(\ell^{2})$, the update cost simplifies to $O((d_{x}+d_{y})\ell)$. Given that there are $\log{R}+1$ levels to update, the total cost of hDS-COD is $O\left(\frac{d_{x}+d_{y}}{\varepsilon}\log{R}\right)$.” (the last paragraph of Appendix C in (Xian et al., 2025)). This indicates that only under the assumption $d_{x}+d_{y}=\Omega(\ell^{2})$ can the $\ell^{3}$ term in complexity $O((d_{x}+d_{y})\ell+\ell^{3})$ be ignored. Yet, as $\varepsilon$ converges to 0, according to Theorem 1 in (Xian et al., 2025), $\ell$ is set to $\ell=1/\varepsilon$, consequently it is often the case that $d_{x}+d_{y}=O(\ell^{2})$ or even $d_{x}+d_{y}=O(\ell)$. Under such conditions, substituting this into complexity $O((d_{x}+d_{y})\ell+\ell^{3})$ means the $\ell^{3}$ term can no longer be neglected, causing the complexity to degrade to $O((d_{x}+d_{y})^{3})$, matching the per-update time complexity of ML-SO-COD under the same condition of $d_{x}+d_{y}=O(\ell)$. Therefore, when higher approximation quality is required, neither ML-SO-COD nor hDS-COD achieves optimal update time complexity.
Additionally, Section 4.3 of (Yao et al., 2025) proposes Adaptive DS-COD (aDS-COD) in Algorithm 6, which optimizes both time and space complexity based on hDS-COD. However, it is explicitly pointed out that this algorithm is only suitable for streams that “have a large R but little fluctuation within the window”, and not for the more strictly defined version in Problem 4, which requires handling arbitrary fluctuations within the window. Thus, for fairness in comparison, we do not include aDS-COD in Table 1 as a competitor to our AeroSketch framework. Nonetheless, we find that our AeroSketch framework can replace the DS-COD component in aDS-COD, further optimizing its time complexity under conditions where $\varepsilon$ is small. We will demonstrate the experimental results of this optimization in the Appendix I.
@@ 2.2.1. Power Iteration
1 $\bm{x}^{(0)}\leftarrow\mathcal{N}(\bm{0},\bm{I}_{\ell})$, $\bm{x}^{(0)}\leftarrow\frac{\bm{x}^{(0)}}{\left\|\bm{x}^{(0)}\right\|_{2}}$
2
3for $i=1$ to $k$ do
4 $\bm{x}^{(i)}\leftarrow\bm{A}^{\top}\bm{A}\bm{x}^{(i-1)}$, $\bm{x}^{(i)}\leftarrow\frac{\bm{x}^{(i)}}{\left\|\bm{x}^{(i)}\right\|_{2}}$
5
[…]
7 return $\hat{\sigma_{1}}^{2},\bm{v}_{1}$
@@ 2.2.2. Simultaneous Iteration
Output: $\bm{Z}\in\mathbb{R}^{d\times k}$: the matrix of the approximate top-$k$ singular vectors; $\bm{\Sigma}\in\mathbb{R}^{k\times k}$: the diagonal matrix of the approximate top-$k$ singular values.
1
[…]
3 $\bm{K}\leftarrow\left(\bm{A}\bm{A}^{\top}\right)^{q}\bm{A}\bm{\Pi}$ // $\bm{K}\in\mathbb{R}^{d\times k}$, takes $\Theta\left(qd\ell k\right)$ time
4 $\bm{Q},\bm{R}\leftarrow\texttt{QR}(\bm{K})$ // $\bm{Q}\in\mathbb{R}^{d\times k}$, takes $\Theta\left(dk^{2}\right)$ time
5 $\bm{M}\leftarrow\bm{Q}^{\top}\bm{A}\bm{A}^{\top}\bm{Q}$ // $\bm{M}\in\mathbb{R}^{k\times k}$, takes $O(d\ell k)$ time
6 $\bm{U},\hat{\bm{\Sigma}}^{2},\bm{U}^{\top}]\leftarrow\texttt{SVD}(\bm{M})$ // takes $O(k^{3})$ time
@@
The comments in Algorithm 2 analyze the time cost of each line. The time complexity of Algorithm 2 is dominated by line 2, which takes $\Theta(qd\ell k)$ time, where $q=\Theta\left(\frac{\log\ell}{\varepsilon_{\text{SI}}}\right)$. If $\varepsilon_{\text{SI}}$ is a constant, then the total time complexity of Algorithm 2 becomes $\Theta(d\ell k\log\ell)$.
@@ 3.1.1. High-level idea
If $\bm{Z}_{i}$ is formed by exact singular vectors of ${\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}$ as in Fast-DS-FD, then Eq. (1) is implicitly zero. However, if the $\bm{Z}_{i}$’s are approximate singular vectors, the restoring norm error becomes explicitly nonzero and can accumulate as the update proceeds, as shown by the blue line in Figure 1 (See Appendix B for a detailed derivation). It is also difficult to derive an upper bound for this error. Fortunately, we found that by redesigning the saved snapshot ($\bm{Z}_{i}$ and $\bm{Z}_{i}^{\top}{\bm{C}^{\prime}_{i}}^{\top}{\bm{C}^{\prime}_{i}}$ at line 15 of Algorithm 4) and the restoring method of $\bm{B}_{\text{W}}$ involving the snapshots (Algorithm 5), this cumulative error can be eliminated. Furthermore, other execution procedures and parameter selections must be meticulously configured to bound the error, space/communication costs, and time complexity. Through analysis and experiments, we find that the restoring norm error becomes zero (as shown by the black line in Figure 1) and the algorithm’s error now stems solely from the FD reduce operation and the residual matrix $\bm{C}_{T-N}$ at time TNT-N, both of which are more tractable for error analysis. Additionally, we discovered that the auxiliary sketch and queue in Fast-DS-FD can be removed, saving their space and time overhead.
@@
1 $\ell\leftarrow\lceil\frac{2}{\varepsilon}\rceil$, $\bm{C}_{0}\leftarrow\bm{0}_{2\ell\times d}$
2 $\mathcal{S}\leftarrow\mathsf{queue}.\textsc{Initialize()}$ // Queue of snapshots
@@ 3.1.2. Data Structures
1
2Insert $\bm{a}_{i}$ into an empty row of $\bm{C}_{i-1}$ to get $\tilde{\bm{C}}_{i}$
3
4if $\text{rows}(\tilde{\bm{C}}_{i})==2\ell$ then // $\tilde{\bm{C}}_{i}$ has no zero rows
5 $\bm{U}_{i},\bm{\Sigma}_{i},\bm{V}_{i}^{\top}]\leftarrow\mathtt{SVD}(\tilde{\bm{C}}_{i})$ // $\bm{\Sigma}_{i}=\text{diag}\left(\sigma_{1},\sigma_{2}\dots,\sigma_{\min\left(2\ell,d\right)}\right)$
6 $\bm{C}_{i}^{\prime}\leftarrow\sqrt{\max\left(\bm{\Sigma}_{i}^{2}-\sigma_{\ell}^{2}\bm{I},\bm{0}\right)}\,\bm{V}_{i}^{\top}$ // element-wise max
7
8else
9 ${\bm{C}}_{i}^{\prime}\leftarrow\tilde{\bm{C}}_{i}$
10
11
[…]
13if $\hat{\sigma}_{1}^{2}\geq\frac{\theta}{2}$ then
14 for $j=1$ to $\lceil\log_{2}\min\left(\ell,d\right)\rceil$ do
15    $k\leftarrow\min(2^{j},\ell,d)$
16    $\bm{Z}_{i}^{\prime},\hat{\bm{\Sigma}}]\leftarrow\mathtt{simul\_iter}({\bm{C}^{\prime}_{i}}^{\top},k,\varepsilon_{\text{SI}}=0.4)$ // $\hat{\bm{\Sigma}}=\text{diag}\left(\hat{\sigma}_{1},\hat{\sigma}_{2}\dots,\hat{\sigma}_{k}\right)$
17    if $\hat{\bm{\sigma}}^{2}_{k}<\theta$ then
18       $\xi\leftarrow\max_{\xi}\{\hat{\bm{\sigma}}_{\xi}^{2}\geq\theta\}$
19       $\bm{Z}_{i}\leftarrow{\bm{Z}_{i}^{\prime}}[:,1:\xi\ ]$
20       $\mathcal{S}.\textsc{append}((\bm{Z}_{i},\bm{Z}_{i}^{\top}{\bm{C}^{\prime}_{i}}^{\top}{\bm{C}^{\prime}_{i}}),s=\mathcal{S}[-1].t+1,t=i)$ // $\mathcal{S}[-1]$ is the tail element in queue $\mathcal{S}$
21       $\bm{C}_{i}\leftarrow{\bm{C}^{\prime}_{i}}-{\bm{C}^{\prime}_{i}}\bm{Z}_{i}\bm{Z}_{i}^{\top}$
22         break
23
24
25 $\bm{C}_{i}\leftarrow\bm{C}_{i}^{\prime}$
26
27else
28 $\bm{C}_{i}\leftarrow\bm{C}_{i}^{\prime}$
29
@@ 3.1.3. Update Algorithm
Algorithm 4 shows how AeroSketch updates the sketch matrix $\bm{C}_{i}$ when a new row vector $\bm{a}_{i}$ arrives at timestamp $i$. First, we insert $\bm{a}_{i}$ into an empty row of $\bm{C}_{i-1}$ to form $\tilde{\bm{C}}_{i}$ (line 1). If the number of rows in $\tilde{\bm{C}}_{i}$ reaches $2\ell$, we apply lines 3-4, the FD reduction described in Section 2.3. Otherwise, we simply set $\bm{C}_{i}^{\prime}$ to $\tilde{\bm{C}}_{i}$ (line 6). Next, we perform Power Iteration on matrix $\bm{C}_{i}^{\prime}$ to estimate the squared top singular value $\hat{\sigma}_{1}^{2}$ (line 7). If it exceeds half the dump threshold, i.e., $\hat{\sigma}_{1}^{2}>\theta/2$ (line 8), we apply Simultaneous Iteration to $\bm{C}_{i}^{\prime}$ to compute the approximate squared top-$k$ singular values greater than $\theta$ along with the corresponding singular vectors. Thus, instead of recomputing the exact singular vectors, we quickly approximate them in a few iterations.
@@
[…]
2 forall $((\bm{Z}_{j},\bm{Z}_{j}^{\top}{\bm{C}_{j}^{\prime}}^{\top}\bm{C}^{\prime}_{j}),s,t=j)$ of $\mathcal{S}$ do
3 if $lb<t=j\leq ub$ then
4    $\bm{B}\leftarrow\bm{B}+\bm{Z}_{j}{\bm{Z}_{j}}^{\top}{\bm{C}^{\prime}_{j}}^{\top}\bm{C}^{\prime}_{j}+{\bm{C}^{\prime}_{j}}^{\top}\bm{C}^{\prime}_{j}\bm{Z}_{j}{\bm{Z}_{j}}^{\top}-\bm{Z}_{j}{\bm{Z}_{j}}^{\top}{\bm{C}^{\prime}_{j}}^{\top}\bm{C}^{\prime}_{j}\bm{Z}_{j}\bm{Z}_{j}^{\top}$
5
6
[…]
8 $\hat{\bm{B}}\leftarrow\sqrt{\max\left(\bm{\Lambda}-\bm{I}\lambda_{\ell},\bm{0}\right)}\,\bm{V}^{\top}$

### v2

@@ 1. Introduction
Many types of real-world data—such as computer network traffic (Xia et al., 2014; Kim et al., 2020), online machine learning (Kuzborskij et al., 2019; Chen et al., 2020; Wen et al., 2025; Chen et al., 2022; Wen et al., 2024), and sensor data (Xie et al., 2023; Aquino et al., 2007)—are continuously generated as vast, high-velocity streams of vectors (Zhou, 2024; Wang et al., 2024; Muthukrishnan, 2005). Storing even a reasonable portion of these streams in memory is impractical. Therefore, it is necessary to design online algorithms that process one vector at a time and immediately discard it, enabling real-time processing with reduced memory usage at the cost of some precision. Such compact representations are known as matrix sketching (Cormode and Yi, 2020; Liberty, 2013; Woodruff, 2014; Desai et al., 2016; Kacham, 2024).
@@ 2.1.2. Matrix Sketching over Sliding Window
Besides the space-optimal Fast-DS-FD algorithm proposed by Yin et al. in (Yin et al., 2024) as we listed in Table 1, they also mentioned a time-optimized version called probabilistic Fast-DS-FD in Theorem 3.1, claiming that the per-update time complexity is $O(d\ell)$, by setting $\ell\leftarrow\min\left(\lceil\frac{1}{\varepsilon}\rceil,d\right)$, so that the time complexity becomes $O\left(\frac{d}{\varepsilon}\right)$. Unfortunately, probabilistic Fast-DS-FD lacks a concrete implementation and therefore remains a potential direction for future work. It also lacks strict proofs of error bounds and complexities. Important parameters are not defined explicitly, such as how to set the number of iterations for the RandNLA algorithm to meet the error bound. Therefore, we did not list it in Table 1 as a competitor to our AeroSketch framework. Additionally, we found that Lemma 1 in (Yin et al., 2024), which involves the key operation of dumping and restoring snapshots and the associated correctness proof, holds only under the premise that $\bm{v}_{j}$ is an exact singular vector. However, the probabilistic randomized algorithm used by probabilistic Fast-DS-FD to optimize time can obtain only an approximate estimate of $\bm{v}_{j}$. Therefore, directly replacing the SVD in Fast-DS-FD with RandNLA will introduce cumulative restoring norm errors, as shown in Figure 1. Designing the dumping and restoring snapshots operation and correctness proof under this premise is far from straightforward. In this paper, our AeroSketch framework will overcome these limitations.
@@ 2.2.1. Power Iteration
$\bm{x}^{(0)}\leftarrow\mathcal{N}(\bm{0},\bm{I}_{\ell})$, $\bm{x}^{(0)}\leftarrow\frac{\bm{x}^{(0)}}{\left\|\bm{x}^{(0)}\right\|_{2}}$
for $i=1$ to $k$ do
 $\bm{x}^{(i)}\leftarrow\bm{A}^{\top}\bm{A}\bm{x}^{(i-1)}$, $\bm{x}^{(i)}\leftarrow\frac{\bm{x}^{(i)}}{\left\|\bm{x}^{(i)}\right\|_{2}}$
end for
$\hat{\sigma_{1}}^{2}\leftarrow\left(\bm{x}^{(k)}\right)^{\top}\bm{A}^{\top}\bm{A}\bm{x}^{(k)},\bm{v}_{1}\leftarrow\bm{x}^{(k)}$
return $\hat{\sigma_{1}}^{2},\bm{v}_{1}$
@@ 2.2.2. Simultaneous Iteration
Output: $\bm{Z}\in\mathbb{R}^{d\times k}$: the matrix of the approximate top-$k$ singular vectors; $\hat{\bm{\Sigma}}\in\mathbb{R}^{k\times k}$: the diagonal matrix of the approximate top-$k$ singular values.
$q\leftarrow\Theta\left(\frac{\log d}{\varepsilon_{\text{SI}}}\right),\bm{\Pi}\sim\mathcal{N}(0,1)^{\ell\times k}$
$\bm{K}\leftarrow\left(\bm{A}\bm{A}^{\top}\right)^{q}\bm{A}\bm{\Pi}$ // $\bm{K}\in\mathbb{R}^{d\times k}$, takes $\Theta\left(qd\ell k\right)$ time
$\bm{Q},\bm{R}\leftarrow\texttt{QR}(\bm{K})$ // $\bm{Q}\in\mathbb{R}^{d\times k}$, takes $\Theta\left(dk^{2}\right)$ time
$\bm{M}\leftarrow\bm{Q}^{\top}\bm{A}\bm{A}^{\top}\bm{Q}$ // $\bm{M}\in\mathbb{R}^{k\times k}$, takes $O(d\ell k)$ time
$\bm{U},\hat{\bm{\Sigma}}^{2},\bm{U}^{\top}]\leftarrow\texttt{SVD}(\bm{M})$ // takes $O(k^{3})$ time
@@
The comments in Algorithm 2 analyze the time cost of each line. The time complexity of Algorithm 2 is dominated by line 2, which takes $\Theta(qd\ell k)$ time, where $q=\Theta\left(\frac{\log d}{\varepsilon_{\text{SI}}}\right)$. If $\varepsilon_{\text{SI}}$ is a constant, then the total time complexity of Algorithm 2 becomes $\Theta(d\ell k\log d)$.
@@ 3.1.1. High-level idea
If $\bm{Z}_{i}$’s are formed by exact singular vectors of ${\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}$ as in Fast-DS-FD, then Eq. (1) is implicitly zero. However, if the $\bm{Z}_{i}$’s are approximate singular vectors, the restoring norm error becomes explicitly nonzero and can accumulate as the update proceeds, as shown by the blue line in Figure 1 (See Appendix B for a detailed derivation). It is also difficult to derive an upper bound for this error. Fortunately, we found that by redesigning the saved snapshot ($\bm{Z}_{i}$ and $\bm{Z}_{i}^{\top}{\bm{C}^{\prime}_{i}}^{\top}{\bm{C}^{\prime}_{i}}$ at line 15 of Algorithm 4) and the restoring method of $\bm{B}_{\text{W}}$ involving the snapshots (Algorithm 5), this cumulative error can be eliminated. Furthermore, other execution procedures and parameter selections must be meticulously configured to bound the error, space/communication costs, and time complexity. Through analysis and experiments, we find that the restoring norm error becomes zero (as shown by the black line in Figure 1) and the algorithm’s error now stems solely from the FD reduce operation and the residual matrix $\bm{C}_{T-N}$ at time TNT-N, both of which are more tractable for error analysis. Additionally, we discovered that the auxiliary sketch and queue in Fast-DS-FD can be removed, saving their space and time overhead.
@@
$\ell\leftarrow\lceil\frac{2}{\varepsilon}\rceil$, $\bm{C}_{0}\leftarrow\bm{0}_{2\ell\times d}$
$\mathcal{S}\leftarrow\mathsf{queue}.\textsc{Initialize()}$ // Queue of snapshots
@@ 3.1.2. Data Structures
Insert $\bm{a}_{i}$ into an empty row of $\bm{C}_{i-1}$ to get $\tilde{\bm{C}}_{i}$
if $\text{rows}(\tilde{\bm{C}}_{i})==2\ell$ then // $\tilde{\bm{C}}_{i}$ has no zero rows
 $\bm{U}_{i},\bm{\Sigma}_{i},\bm{V}_{i}^{\top}]\leftarrow\mathtt{SVD}(\tilde{\bm{C}}_{i})$ // $\bm{\Sigma}_{i}=\text{diag}\left(\sigma_{1},\sigma_{2}\dots,\sigma_{\min\left(2\ell,d\right)}\right)$
 $\bm{C}_{i}^{\prime}\leftarrow\sqrt{\max\left(\bm{\Sigma}_{i}^{2}-\sigma_{\ell}^{2}\bm{I},\bm{0}\right)}\,\bm{V}_{i}^{\top}$ // element-wise max
else
 ${\bm{C}}_{i}^{\prime}\leftarrow\tilde{\bm{C}}_{i}$
end if
$\hat{\sigma}_{1}^{2},\hat{\bm{v}}_{1}\leftarrow\mathtt{power\_iteration}({\bm{C}^{\prime}_{i}}^{\top},k=\lceil\log_{2}d\rceil+1)$
if $\hat{\sigma}_{1}^{2}\geq\frac{\theta}{2}$ then
 for $j=1$ to $\lceil\log_{2}\min\left(\ell,d\right)\rceil$ do
    $k\leftarrow\min(2^{j},\ell,d)$
    $\bm{Z}_{i}^{\prime},\hat{\bm{\Sigma}}]\leftarrow\mathtt{simul\_iter}({\bm{C}^{\prime}_{i}}^{\top},k,\varepsilon_{\text{SI}}=0.4)$ // $\hat{\bm{\Sigma}}=\text{diag}\left(\hat{\sigma}_{1},\hat{\sigma}_{2}\dots,\hat{\sigma}_{k}\right)$
    if $\hat{\bm{\sigma}}^{2}_{k}<\theta$ then
       $\xi\leftarrow\max_{\xi}\{\hat{\bm{\sigma}}_{\xi}^{2}\geq\theta\}$
       $\bm{Z}_{i}\leftarrow{\bm{Z}_{i}^{\prime}}[:,1:\xi\ ]$
       $\mathcal{S}.\textsc{append}((\bm{Z}_{i},\bm{Z}_{i}^{\top}{\bm{C}^{\prime}_{i}}^{\top}{\bm{C}^{\prime}_{i}}),s=\mathcal{S}[-1].t+1,t=i)$ // $\mathcal{S}[-1]$ is the tail element in queue $\mathcal{S}$
       $\bm{C}_{i}\leftarrow{\bm{C}^{\prime}_{i}}-{\bm{C}^{\prime}_{i}}\bm{Z}_{i}\bm{Z}_{i}^{\top}$
         break
      end if
   end for
 $\bm{C}_{i}\leftarrow\bm{C}_{i}^{\prime}$
else
 $\bm{C}_{i}\leftarrow\bm{C}_{i}^{\prime}$
end if
@@ 3.1.3. Update Algorithm
Algorithm 4 shows how AeroSketch updates the sketch matrix $\bm{C}_{i}$ when a new row vector $\bm{a}_{i}$ arrives at timestamp $i$. First, we insert $\bm{a}_{i}$ into an empty row of $\bm{C}_{i-1}$ to form $\tilde{\bm{C}}_{i}$ (line 1). If the number of rows in $\tilde{\bm{C}}_{i}$ reaches $2\ell$, we apply lines 3-4, the FD reduction described in Section 2.3 to form $\bm{C}_{i}^{\prime}$. Otherwise, we simply set $\bm{C}_{i}^{\prime}$ to $\tilde{\bm{C}}_{i}$ (line 6). Next, we perform Power Iteration on matrix $\bm{C}_{i}^{\prime}$ to estimate the squared top singular value $\hat{\sigma}_{1}^{2}$ (line 7). If it exceeds half the dump threshold, i.e., $\hat{\sigma}_{1}^{2}>\theta/2$ (line 8), we apply Simultaneous Iteration to $\bm{C}_{i}^{\prime}$ to compute the approximate squared top-$k$ singular values greater than $\theta$ along with the corresponding singular vectors. Thus, instead of recomputing the exact singular vectors, we quickly approximate them in a few iterations.
@@
$\bm{B}\leftarrow\bm{C}_{ub}^{\top}\bm{C}_{ub}$
forall $((\bm{Z}_{j},\bm{Z}_{j}^{\top}{\bm{C}_{j}^{\prime}}^{\top}\bm{C}^{\prime}_{j}),s,t=j)$ of $\mathcal{S}$ do
 if $lb<t=j\leq ub$ then
    $\bm{B}\leftarrow\bm{B}+\bm{Z}_{j}{\bm{Z}_{j}}^{\top}{\bm{C}^{\prime}_{j}}^{\top}\bm{C}^{\prime}_{j}+{\bm{C}^{\prime}_{j}}^{\top}\bm{C}^{\prime}_{j}\bm{Z}_{j}{\bm{Z}_{j}}^{\top}-\bm{Z}_{j}{\bm{Z}_{j}}^{\top}{\bm{C}^{\prime}_{j}}^{\top}\bm{C}^{\prime}_{j}\bm{Z}_{j}\bm{Z}_{j}^{\top}$
   end if
end forall
$\bm{V},\bm{\Lambda},\bm{V}^{\top}]\leftarrow\mathtt{eigen\_decomposition}(\bm{B})$ // $\bm{\Lambda}=\text{diag}\left(\lambda_{1},\lambda_{2}\dots,\lambda_{d}\right),\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{d}$
$\hat{\bm{B}}\leftarrow\sqrt{\max\left(\bm{\Lambda}-\bm{I}\lambda_{\ell},\bm{0}\right)}\,\bm{V}^{\top}$